# Dysca: A Dynamic and Scalable Benchmark for Evaluating Perception Ability of LVLMs

### v2

@@ Abstract
Currently many benchmarks have been proposed to evaluate the perception ability of the Large Vision-Language Models (LVLMs). However, most benchmarks conduct questions by selecting images from existing datasets, resulting in the potential data leakage. Besides, these benchmarks merely focus on evaluating LVLMs on the realistic style images and clean scenarios, leaving the multi-stylized images and noisy scenarios unexplored. In response to these challenges, we propose a dynamic and scalable benchmark named Dysca for evaluating LVLMs by leveraging synthesis images. Specifically, we leverage Stable Diffusion and design a rule-based method to dynamically generate novel images, questions and the corresponding answers. We consider 51 kinds of image styles and evaluate the perception capability in 20 subtasks. Moreover, we conduct evaluations under 4 scenarios (i.e., Clean, Corruption, Print Attacking and Adversarial Attacking) and 3 question types (i.e., Multi-choices, True-or-false and Free-form). Thanks to the generative paradigm, Dysca serves as a scalable benchmark for easily adding new subtasks and scenarios. A total of 8 advanced open-source LVLMs with 10 checkpoints are evaluated on Dysca, revealing the drawbacks of current LVLMs. The benchmark is released in https://github.com/Benchmark-Dysca/Dysca.
footnotetext: Equal contribution.
Corresponding Author: zhangjie@ict.ac.cn
@@ 1 Introduction
Recent years have witnessed the great success of the Large Vision-Language Models (LVLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. These models leverage the powerful Large Language Models (LLMs) [11, 12, 13, 14, 15] as their brain and incorporate the state-of-the-art visual encoders [16, 17, 18] as their eyes. Thanks to the alignment of visual feature with textual space and the development of visual instruction tuning techniques [4], LVLMs showcase the impressive capability in terms of visual scene comprehension and multimodal instruction-following.
@@
In order to comprehensively evaluate the capabilities of LVLMs, many benchmarks have been purposed [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], where we categorize the current benchmarks into three types [25]. The first type is the classical benchmarks, such as COCO Caption [30] and VQA [19, 31, 32]. Although these benchmarks provide high-quality evaluation data, they also have notable limitations. On the one hand, they are inadequate for measuring the fine-grained capabilities of current LVLMs, offering the limited insightful feedback for the future improvement. On the other hand, since these classical benchmarks have been available as the open-source test data for a long time, it is hard to prevent the data leakage problem. The second type of benchmarks evaluate the LVLMs through a subjective manner [28, 33]. Although the benchmarks reveal the insightful drawbacks of current models, their data scale is limited (i.e., less than 200 annotations) and they require manual evaluation by experts. The third type is built for objectively evaluating current LVLMs and the comparison between them are shown in Tab. 1. They provide an objective and automatic evaluation manner, giving the fine-grained evaluation for the LVLMs. However, these benchmarks conduct Vision-language QAs by selecting images from existing dataset. Although they claim that the questions are re-annotated, the previous work [29] has demonstrated that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs. Besides, most benchmarks focus on evaluating LVLMs in the realistic images and clean scenarios, leaving the multi-stylized images and noisy scenarios unexplored. While some works like MMCBench [34] and Typographic Dataset [35] have investigated the robustness of LVLMs with corrupted and print-attacked images, respectively, they have not explored the effect of these noisy images on various perceptual tasks.
@@
Figure 1: Overview of the automatic pipeline for generating Vision-language QAs, cleaning Vision-language QAs and evaluating LVLMs. (a) We first constructs prompts in terms of content, style and background, leveraging the Text-to-Image (T2I) diffusion model (e.g., SDXL [36]) to synthesis images to be asked. Then based on the scenarios and the question type, we post-process the synthesis images and generate the specific textual questions, respectively. (b) We further filter out low quality Vision-language QAs by utilizing trained models to form the final Dysca. (c) Finally, we evaluate LVLMs on our Dysca and feedback the fine-grained evaluation results.
@@
In this paper, aiming to address these challenges above, we propose Dysca which is a dynamic and scalable benchmark for evaluating the perception ability of LVLMs via various subtasks and scenarios. Inspired by the prior evaluation works for LLMs [37], we investigate on whether we could leverage the large-scale synthesized images for evaluating LVLMs. We display the overview of our pipeline in Fig. 1. Specifically, we leverage Stable Diffusion and design a rule-based method to dynamically generate novel images, questions and corresponding answers. We decouple the prompt into 4 part, i.e., attribute, foreground, style and background, and design pre-defined templates to dynamically generate prompts, as displayed in Fig. 3. Then we utilize the state-of-the-art text-to-image diffusion models (e.g., SDXL [36]) to generate the corresponding images. Since we already know the main information of the images through prompts, we easily generate question-answer textual pairs by the rule-based method. After that, in order to obtain the high quality Vision-language QAs, we employ CLIP [16] to perform data cleaning on the generated Vision-language QA pairs. Dysca focuses on assessing the fine-grained perception abilities, including recognizing human, animal, object, landmark, etc. Dysca evaluates LVLMs with 20 perceptual subtasks, containing a total number of 51 different artistic styles. Besides, to evaluate the robustness of the models across different scenarios and question types, we construct 4 testing scenarios (clean, corruption, print attacking and adversarial attacking) and 3 question types (multi-choices, true-or-false and free-form questions). In the end, Dysca consists of 617K Vision-language QA pairs ($\times$20 larger than MM-BigBench [38] and $\times$25 larger than Seed-Bench2 [24] as shown in Tab. 1). Thanks to the generative paradigm, Dysca achieves the scalable benchmark to new subtasks and scenarios and dynamically generate unlimited Vision-language QAs for evaluation.
@@
  • Multi-grained Perceptual Subtasks and Multi-scenarios: Dysca evaluates the 20 perceptual subtasks performance of 8 mainstream LVLMs with 10 checkpoints under 4 image scenarios (i.e., clean, corruption, print attacking and adversarial attacking) and 3 question types (i.e., multi-choices, true-or-false and free-form questions).
@@
Collecting from Existing Datasets Question Type
@@
LLaVA-Bench [4] 0.15K - Free-form
MME [25] 2.3K 10 True-or-false
LVLM-eHub [21] - 3 Free-form
tiny-LVLM-eHub [22] 2.1K 3 Free-form
SEED-Bench [23] 19K 8 Multi-choices
MMBench [39] 2.9K 12 Multi-choices
TouchStone [26] 0.9K 10 Free-form
REFORM-EVAL [40] 50K 7 Multi-choices
MM-BigBench [38] 30K 6 Multi-choices
MM-VET [27] 0.2K 4 Free-form
MLLM-Bench [41] 0.42K 7 Free-form
SEED-Bench2 [24] 24K 10 Multi-choices
BenchLMM [42] 2.4K 15 Free-form
JourneyDB [43] 5.4K 2
@@ 2.1 Large Vision-Language Models
The landscape of Large Vision-Language Models (LVLMs) has been significantly shaped by the pioneering success of Large Language Models (LLMs) such as GPTs [44, 45, 46] and LLaMA [13], catalyzing advancements in multimodal content understanding and generation [47], including intricate tasks like image-text comprehension. At the forefront of these developments, BLIP-2 [1] introduces a lightweight Q-Former [1] that facilitates alignment between textual and visual representations through a cross-attention mechanism [1]. InstructBLIP [3] takes a step further by incorporating textual instructions into the Q-Former, which significantly improves performance. LLAVA [4] employs GPT-4 [14] to transform data into multimodal instruction-following data and uses CLIP [16] and LLAMA [13] for fine-tuning instructions, achieving advanced performance. LLAVA-1.5 [48] extends this paradigm by integrating MLP projection and introducing academic task-specific Vision-language QA data. Recently, models like Otter [5], MiniGPT-4 [2], Qwen-VL-Chat [49] and XComposer-VL [7] further unleash the cross-modal understanding capabilities of LVLMs.
@@ 2.2 Benchmarks for LVLMs
The great progress of LVLMs triggers the development of benchmarks for evaluating these models, where we divide previous benchmarks into three categories. The first type is the classical benchmarks which focuses on evaluating LVLMs abilities via image caption [50] and VQA [51, 19]. However, these benchmarks cannot provide the fine-grained feedback on how to improve the models. Besides, since these benchmarks have been the public resources for a long time, it is hardly to guarantee that the LVLMs have not use them for training. The second type subjectively evaluates LVLMs by experts [28, 33]. Although these benchmarks reveal the insightful feedback of the LVLMs, their scale is limited (i.e., less than 200 annotations). The subjective manner also makes the evaluation expensive and hardly to expand the scale of the benchmarks.
The third type [4, 25, 21, 22, 23, 24, 39, 26, 40, 38, 27, 41, 42, 29, 52] focuses on evaluating LVLMs in an objective and large-scaled manner, where we list the detailed information of them in the Tab. 1. Some of them have been adopted by the community [53] as the standard benchmarks for evaluating LVLMs [12, 5, 7], like MME [25] and MMBench [39]. These benchmarks evaluate models through the objective answer types and most of them leverage the automatic annotation and evaluation manner for revealing the fine-grained drawbacks of current LVLMs. However, the previous benchmarks primarily concentrate on evaluating LVLMs using realistic images and clean scenario, leaving multi-stylized images and noisy scenarios unexplored. Moreover, many of them conduct QA by selecting images from publicly available datasets (e.g., [50, 54]). While they state that the questions have been re-annotated, they cannot guarantee that the LVLMs have not seen the image during training stage. The previous work [29] has proved that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs. One possible way to solve data leakage is using novel but synthesis images, where JourneyDB [43] is the first work aiming to leverage synthesis images to evaluate current LVLMs. The prompts and the corresponding images are downloaded from Midjourney [55] and ChatGPT [12] is leveraged to label the images. However, JourneyDB is a top-down framework where the number of images is fixed. Besides, the ChatGPT labeling may cause hallucinate annotations, leading to the unreliable evaluation results. Although 40 annotators have involved to clean the data, the data cleaning cost are expensive and it limits the data scale. In contrast, our Dysca serves as the bottom-up framework, allowing for dynamic and scalable generation for both images and evaluation questions. The rule-based question generation method also makes the annotations more accuracy. Besides, Dysca contains 20 subtasks which is more comprehensive than JourneyDB.
@@ 3.1 Overview of Our Pipeline
The overview of our pipeline is shown in Fig. 1, containing data generation, data cleaning and LVLMs evaluation. For the data generation, our Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, where $M$ means "Metadata", $P$ means "Prompt", $I$ means "Image" and $Q$ means "Question-answer pair". We further decouple the metadata $M$ into 4 parts, i.e., "style", "attribute", "foreground" and "background", and the combination of the four parts constitute the image prompts $P$. Then, given the prompt $P$ and the selected scenario, we leverage the Text-to-Image (T2I) diffusion model (e.g., SDXL [36]) to synthesis image $I$ and add the specific perturbation to the image $I$. After that, since the prompt already includes the question angle and the corresponding answer, we construct a rule-based approach to generate the $Q$. Three types of questions are considered, i.e., multi-choice, true-or-false and free-form. Multi-choice and true-or-false questions utilize a closed-ended manner to assess LVLMs, while free-form questions employ an open-ended manner through image captioning for evaluation. For the data cleaning, considering that the T2I diffusion model may generate unsuccessful outcomes, we then use CLIP [16] and PP-OCRv3 [56] to automatically clean the whole dataset to obtain the final Dysca. Finally, we evaluate 8 open-source LVLMs with 10 checkpoints on our proposed Dysca.

### v3

@@ Abstract
Currently many benchmarks have been proposed to evaluate the perception ability of the Large Vision-Language Models (LVLMs). However, most benchmarks conduct questions by selecting images from existing datasets, resulting in the potential data leakage. Besides, these benchmarks merely focus on evaluating LVLMs on the realistic style images and clean scenarios, leaving the multi-stylized images and noisy scenarios unexplored. In response to these challenges, we propose a dynamic and scalable benchmark named Dysca for evaluating LVLMs by leveraging synthesis images. Specifically, we leverage Stable Diffusion and design a rule-based method to dynamically generate novel images, questions and the corresponding answers. We consider 51 kinds of image styles and evaluate the perception capability in 20 subtasks. Moreover, we conduct evaluations under 4 scenarios (i.e., Clean, Corruption, Print Attacking and Adversarial Attacking) and 3 question types (i.e., Multi-choices, True-or-false and Free-form). Thanks to the generative paradigm, Dysca serves as a scalable benchmark for easily adding new subtasks and scenarios. A total of 24 advanced open-source LVLMs and 2 close-source LVLMs are evaluated on Dysca, revealing the drawbacks of current LVLMs. The benchmark is released at https://github.com/Robin-WZQ/Dysca.
@@ 1 Introduction
Recent years have witnessed the great success of the Large Vision-Language Models (LVLMs) (Li et al., 2023d; Zhu et al., 2023; Dai et al., 2023; Liu et al., 2023b; Li et al., 2023a; Chen et al., 2023b; Zhang et al., 2023; Su et al., 2023; Gong et al., 2023; Sun et al., 2023b). These models leverage the powerful Large Language Models (LLMs) (Chung et al., 2022; OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023; FastChat, 2023) as their brain and incorporate the state-of-the-art visual encoders (Radford et al., 2021; Fang et al., 2023; Dosovitskiy et al., 2020) as their eyes. Thanks to the alignment of visual feature with textual space and the development of visual instruction tuning techniques (Liu et al., 2023b), LVLMs showcase the impressive capability in terms of visual scene comprehension and multimodal instruction-following.
@@
In order to comprehensively evaluate the capabilities of LVLMs, many benchmarks have been purposed (Antol et al., 2015a; Singh et al., 2019; Xu et al., 2023; Shao et al., 2023; Li et al., 2023c; b; Fu et al., 2023; Bai et al., 2023b; Yu et al., 2023; Yang et al., 2023b; Chen et al., 2024), where we categorize the current benchmarks into three types (Fu et al., 2023). The first type is the classical benchmarks, such as COCO Caption (Chen et al., 2015) and VQA (Antol et al., 2015a; Goyal et al., 2017; Marino et al., 2019). Although these benchmarks provide high-quality evaluation data, they also have notable limitations. On the one hand, they are inadequate for measuring the fine-grained capabilities of current LVLMs, offering the limited insightful feedback for the future improvement. On the other hand, since these classical benchmarks have been available as the open-source test data for a long time, it is hard to prevent the data leakage problem. The second type of benchmarks evaluate the LVLMs through a subjective manner (Yang et al., 2023b; Wu et al., 2023). Although the benchmarks reveal the insightful drawbacks of current models, their data scale is limited (i.e., less than 200 annotations) and they require manual evaluation by experts. The third type is built for objectively evaluating current LVLMs and the comparison between them are shown in Tab. 1. They provide an objective and automatic evaluation manner, giving the fine-grained evaluation for the LVLMs. However, these benchmarks conduct Vision-language QAs by selecting images from existing dataset and annotate the textual questions. Although they claim that the questions are re-annotated, the previous work (Chen et al., 2024) has demonstrated that these benchmarks have Models unintentionally leaked into the training data of LLMs and LVLMs. Besides, most benchmarks focus on evaluating LVLMs in the realistic images and clean scenarios, leaving the multi-stylized images and noisy scenarios unexplored. While some works like MMCBench (Zhang et al., 2024b) and Typographic Dataset (Cheng et al., 2024) have investigated the robustness of LVLMs with corrupted and print-attacked images, respectively, they have not explored the effect of these noisy images on various perceptual tasks.
@@
Figure 1: Overview of the automatic pipeline for generating Vision-language QAs, cleaning Vision-language QAs and evaluating LVLMs. (a) We first constructs prompts in terms of content, style and background, leveraging the Text-to-Image (T2I) diffusion model (e.g., SDXL (Podell et al., 2023)) to synthesis images to be asked. Then based on the scenarios and the question type, we post-process the synthesis images and generate the specific textual questions, respectively. (b) We further filter out low quality Vision-language QAs by utilizing trained models to form the final Dysca. (c) Finally, we evaluate LVLMs on our Dysca and feedback the fine-grained evaluation results.
@@
In this paper, aiming to address these challenges above, we propose Dysca which is a dynamic and scalable benchmark for evaluating the perception ability of LVLMs via various subtasks and scenarios. Inspired by the prior evaluation works for LLMs (Liang et al., 2023), we investigate on whether we could leverage the large-scale synthesized images for evaluating LVLMs. We display the overview of our pipeline in Fig. 1. Specifically, we leverage Stable Diffusion and design a rule-based method to dynamically generate novel images, questions and corresponding answers. We decouple the prompt into 4 part, i.e., attribute, foreground, style and background, and design pre-defined templates to dynamically generate prompts, as displayed in Fig. 3. Then we utilize the state-of-the-art text-to-image diffusion models (e.g., SDXL (Podell et al., 2023)) to generate the corresponding images. Since we already know the main information of the images through prompts, we easily generate question-answer textual pairs by the rule-based method. After that, in order to obtain the high quality Vision-language QAs, we employ CLIP (Radford et al., 2021) to perform data cleaning on the generated Vision-language QA pairs. Dysca focuses on assessing the fine-grained perception abilities, including recognizing human, animal, object, landmark, etc. Dysca evaluates LVLMs with 20 perceptual subtasks, containing a total number of 51 different artistic styles. Besides, to evaluate the robustness of the models across different scenarios and question types, we construct 4 testing scenarios (clean, corruption, print attacking and adversarial attacking) and 3 question types (multi-choices, true-or-false and free-form questions).
Compared to previous works in Tab. 1, we provide an end-to-end process from image to Vision-QA generation. The approach significantly reduces annotation costs compared to manually labeling images (e.g., MME (Fu et al., 2023)) while achieving the correctness for evaluating LVLMs. It also avoids the risk of hallucinate annotations that may occur when using ChatGPT for labeling based on image prompts (e.g., JourneyDB (Sun et al., 2023a)). This novel pipeline enables us to create a benchmark that is easily scalable and adaptable for incorporating new subtasks and scenarios. Thanks to the generative paradigm, Dysca can be customized to meet the specific requirements of the evaluator for testing purposes.
@@
  • Multi-grained Perceptual Subtasks and Multi-scenarios: Dysca aims to testing LVLMs’ performance on diverse styles, 4 image scenarios (i.e., clean, corruption, print attacking and adversarial attacking) and 3 question types (i.e., multi-choices, true-or-false and free-form questions), reporting the 20 perceptual subtasks performance of 26 mainstream LVLMs, including GPT-4o (Ope, ) and Gemini-1.5-Pro (Team et al., 2024).
@@
Novel Images &\&& Novel Questions Question Type
@@
LLaVA-Bench 0.15K - Free-form
MME 2.3K 10 True-or-false
LVLM-eHub - 3 Free-form
tiny-LVLM-eHub 2.1K 3 Free-form
SEED-Bench 19K 8 Multi-choices
MMBench 2.9K 12 Multi-choices
TouchStone 0.9K 10 Free-form
REFORM-EVAL 50K 7 Multi-choices
MM-BigBench 30K 6 Multi-choices
MM-VET 0.2K 4 Free-form
MLLM-Bench 0.42K 7 Free-form
SEED-Bench2 24K 10 Multi-choices
BenchLMM 2.4K 15 Free-form
JourneyDB 5.4K 2
@@ 2.1 Large Vision-Language Models
The landscape of Large Vision-Language Models (LVLMs) has been significantly shaped by the pioneering success of Large Language Models (LLMs) such as GPTs (Radford et al., 2019; Brown et al., 2020; Ouyang et al., 2022) and LLaMA (Touvron et al., 2023), catalyzing advancements in multimodal content understanding and generation (Zhang et al., 2024a), including intricate tasks like image-text comprehension. At the forefront of these developments, BLIP-2 (Li et al., 2023d) introduces a lightweight Q-Former (Li et al., 2023d) that facilitates alignment between textual and visual representations through a cross-attention mechanism (Li et al., 2023d). InstructBLIP (Dai et al., 2023) takes a step further by incorporating textual instructions into the Q-Former, which significantly improves performance. LLAVA (Liu et al., 2023b) employs GPT-4 (OpenAI, 2023) to transform data into multimodal instruction-following data and uses CLIP (Radford et al., 2021) and LLAMA (Touvron et al., 2023) for fine-tuning instructions, achieving advanced performance. LLAVA-1.5 (Liu et al., 2023a) extends this paradigm by integrating MLP projection and introducing academic task-specific Vision-language QA data. Recently, models like Otter (Li et al., 2023a), MiniGPT-4 (Zhu et al., 2023), Qwen-VL-Chat (Bai et al., 2023a) and XComposer-VL (Zhang et al., 2023) further unleash the cross-modal understanding capabilities of LVLMs. Besides, many powerful closed-source LVLMs, including Gemini-1.5-Pro (Team et al., 2024) and GPT-4o (Ope, ), have publicly released their APIs, promoting the development of downstream applications.
@@ 2.2 Benchmarks for LVLMs
The great progress of LVLMs triggers the development of benchmarks for evaluating these models, where we divide previous benchmarks into three categories. The first type is the classical benchmarks which focuses on evaluating LVLMs abilities via image caption (Lin et al., 2014) and VQA (Antol et al., 2015b; a). However, these benchmarks cannot provide the fine-grained feedback on how to improve the models. Besides, since these benchmarks have been the public resources for a long time, it is hardly to guarantee that the LVLMs have not use them for training. The second type subjectively evaluates LVLMs by experts (Yang et al., 2023b; Wu et al., 2023). Although these benchmarks reveal the insightful feedback of the LVLMs, their scale is limited (i.e., less than 200 annotations). The subjective manner also makes the evaluation expensive and hardly to expand the scale.
The third type (Liu et al., 2023b; Fu et al., 2023; Xu et al., 2023; Shao et al., 2023; Li et al., 2023c; b; Liu et al., 2023c; Bai et al., 2023b; Li et al., 2023f; Yang et al., 2023a; Yu et al., 2023; Ge et al., 2023; Cai et al., 2023; Chen et al., 2024; Liu et al., 2024) focuses on evaluating LVLMs in an objective and large-scaled manner, where we list the detailed information of them in the Tab. 1. Some of them have been adopted by the community (Contributors, 2023) as the standard benchmarks for evaluating LVLMs (OpenAI, 2022; Zhang et al., 2023), like MME (Fu et al., 2023) and MMBench (Liu et al., 2023c). These benchmarks evaluate models through the objective answer types and most of them leverage the automatic annotation and evaluation manner for revealing the fine-grained drawbacks of current LVLMs. However, the previous benchmarks primarily concentrate on evaluating LVLMs using realistic images and clean scenario, leaving multi-stylized images and noisy scenarios unexplored. Moreover, many of them conduct QA by selecting images from publicly available datasets (e.g., (Lin et al., 2014; Russakovsky et al., 2014)). While they state that the questions have been re-annotated, they cannot guarantee that the LVLMs have not seen the image during training stage. The previous work (Chen et al., 2024) has proved that these benchmarks have unintentionally leaked into the training data of LLMs and LVLMs. One possible way to solve data leakage is using novel but synthesis images, where JourneyDB (Sun et al., 2023a) is the first work aiming to leverage synthesis images to evaluate current LVLMs. The prompts and the corresponding images are downloaded from Midjourney (mid, ) and ChatGPT (OpenAI, 2022) is leveraged to label the images. However, JourneyDB is a top-down framework where the number of images is fixed. Besides, the ChatGPT labeling may cause hallucinate annotations, leading to the unreliable evaluation results. In contrast, our Dysca serves as the bottom-up framework, allowing for dynamic and scalable generation for both images and evaluation questions and involving more evaluation tasks. The rule-based question generation method also makes the annotations more accuracy.
@@ 3.1 Overview of Our Pipeline
The overview of our pipeline is shown in Fig. 1, containing data generation, data cleaning and LVLMs evaluation. For the data generation, our Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, where $M$ means “Metadata”, $P$ means “Prompt”, $I$ means “Image” and $Q$ means “Question-answer pair”. We further decouple the metadata $M$ into 4 parts, i.e., “style”, “attribute”, “foreground” and “background”, and the combination of the four parts constitute the image prompts $P$. Then, given the prompt $P$ and the selected scenario, we leverage the Text-to-Image (T2I) diffusion model (e.g., SDXL (Podell et al., 2023)) to synthesis image $I$ and add the specific perturbation to the image $I$. After that, since the prompt already includes the question angle and the corresponding answer, we construct a rule-based approach to generate the $Q$. Three types of questions are considered, i.e., multi-choice, true-or-false and free-form. Multi-choice and true-or-false questions utilize a closed-ended manner to assess LVLMs, while free-form questions employ an open-ended manner through image captioning for evaluation. For the data cleaning, considering that the T2I diffusion model may generate unsuccessful outcomes, we then use CLIP (Radford et al., 2021) and PP-OCRv3 (Li et al., 2022) to automatically clean the whole dataset to obtain the final Dysca. Finally, we evaluate 14 open-sourced LVLMs and 2 closed-source LVLMs on our proposed Dysca.
Table 2: Key statistics of Dysca.